following equations respectively:

Demand: P= 1400 - 2Qd

Supply: P = 200 + 10sI

1.1

Calculate the equilibrium price and quantity of shoes.

1.2

Assume that the price of shoes is R700. Use your answer in 1.1 to exp

the resulting situation in the market for shoes, and how equilibrium

be restored without government intervention, ceteris paribus.


1
Expert's answer
2021-11-16T10:41:11-0500

1.1) Market equilibrium occurs when "P_d=P_s":


Equilibrium quantity, QE will be given by:


"1400-2Q_E=200+10Q_E,""1200=12Q_E,""Q_E=100"

Then we can find equilibrium price:



"P_E=1400-2\cdot100=\$1200"



1.2)

At R700,the quantity demanded of shoes by buyers will increase while the quantity supplied by sellers will decrease. This will mean there will be a shortage of shoes created by the excess demand and less supply.

As a result of this shortage, sellers will recognize that they have an opportunity to make higher profits by selling the shoes at a higher price. As a result, the price rises toward the equilibrium level.